# Efficient Fermi-Hubbard model ground-state preparation by coupling to a classical reservoir in the instantaneous-response limit

### v1

@@ IV Conclusion
We show that an instantaneous simplification of the classical reservoir, constrained to its instantaneous-response limit, can serve as an efficient method for ground-state preparation, offering a more feasible approach for near-term quantum computers.
In comparing this work to previous works, we find that the classical reservoir method produces an ansatz expression analogous to the NP ansatz; however, the motivation arises from cooling dynamics rather than the adiabatic theorem, which underpins the HVA. The classical reservoir method exploits total spin symmetry and a closure property of the Lie algebra in the Hubbard hopping terms. This allows the classical reservoir method to prepare a more accurate ground state in the strongly correlated regime, while also reducing implementation costs in the following ways:
@@
    The closure properties of the Lie algebra allow us to use a one-dimensional indexing scheme, eliminating the need for fermionic swap (fSWAP) gates [44] to handle nonadjacent vertical hopping terms in a two-dimensional system. This concept has also been introduced as the excitation-preserving (EP) ansatz in Ref. [35]. However, Ref. [35] reported that the EP ansatz often underperforms compared to the NP ansatz in two-dimensional systems. In contrast, we find that incorporating total spin into the consideration resolves this issue.
Using the classical reservoir method, we find high-accuracy results —infidelities reach the range of $10^{-3}-10^{-6}$ for 8- and 10-site two-dimensional systems, even in the strong correlation region ($U/t{=}8$).
V Acknowledgment
@@ Appendix A Initial state examples
The only requirement for the initial state is to have the target spin value. But, we find it is advantageous to intersperse the doubly occupied sites and the empty sites. Hence, we select the simplest states that fulfill this requirement, which is a product state. We illustrate this with an eight-site example.
@@ Appendix B Numerical optimization details
We use the L-BFGS algorithm from the GPU-supported PyTorch package for optimization [46]. The parameter amplitudes are initialized as random values within the range $[-0.001,0.001]$. Since we lack prior knowledge about the optimal amplitude values, one might consider initializing them all to zero. However, we found that this approach leads to numerical instability. By choosing a small random initialization, we ensure numerical stability while keeping the values small enough to most likely avoid divergence in the optimized energy.
Though our goal is to implement this algorithm on a quantum computer, we are currently using classical computer simulations. For the exponential of the hopping terms, we use an SU(2) identity [47] to implement the hopping terms instead of directly calculating the matrix exponential, which speeds up the computation. This is given by
@@
Appendix C Two leg ladder system
In Fig. C1, we present numerical results for $2\times 4$ and $2\times 5$ ladders in both the weakly and strongly correlated regimes. Note that the ground-state total spin $S$ at $U/t=2$ depends on the number of sites $N$. Specifically, it is a singlet for $N=4n$ (e.g., $N=4,8,12,\dots$) and a triplet for $N=4n+2$ (e.g., $N=6,10,\dots$[32]. In the case of a weakly correlated region ($U/t=2$), when compared on a linear-connectivity quantum machine, no clear pattern emerges. In the $2\times 5$ system, the NP ansatz [33] requires 432 parameters and a circuit depth of 48 to achieve a fidelity of 0.99, while this work requires 475 parameters and a circuit depth of 75 to reach a fidelity of 0.95 (with an relative error of 0.01). In the $2\times 4$ system, however, this work performs better, achieving a fidelity of 0.99 with 150 parameters and a circuit depth of 30, whereas the NP ansatz [33] requires 392 parameters and a circuit depth of 56 to achieve the same fidelity.
Figure C1: Relative errors (solid lines) and infidelities (dashed lines) for the two-leg ladder system with 8 sites and 10 sites. Panel (a) corresponds to $U/t=2$, while panel (b) corresponds to $U/t=8$.
Figure C2: Gradient norms in panels (a) and (b) and relative errors in panels (c) and (d) during optimization over different numbers of layers. The $x$-axis represents the iteration index. The left panels, (a) and (c), correspond to the $2\times 4$ system, while the right panels, (b) and (d), correspond to the $2\times 5$ system.
From the ladder system data, this method offers two key improvements. First, in the strong correlation regime ($U/t=8$), subsequent work on the NP ansatz [35] reported that the largest system investigated, a $3\times 3$ lattice with open boundary conditions, achieved a relative error (defined as $(E-E_{gs})/N$) that mostly saturates around 0.01. In contrast, for the largest system studied here—a $2\times 5$ ladder system—we achieve higher accuracy, with a similarly defined relative error (given by $(E_{gs}-E)/E_{gs}$) of 0.001 when $E_{gs}=-4.15$.
Second, this work demonstrates a potential pathway (employing a one-dimensional indexing scheme) for efficiently simulating multi-dimensional systems that were previously considered unsatisfactory [35]. An intriguing observation arises when comparing the $2\times 4$ ladder system to a two-dimensional periodic cluster. The latter, with increased site connectivity, surprisingly yields better results. This indicates that the classical reservoir approach can be well-suited for simulating realistic three-dimensional lattices with higher connectivity.
Additionally, in Fig. C2, we provide details of the convergence and the number of iterations required as the system size increases. We present results for $U/t=8$ only, as the total spin value of the ground state differs in the $U/t=2$ region for the $2\times 4$ and $2\times 5$ systems, which can affect performance. The results show that optimization can reduce the relative error to below 0.01 relatively quickly, typically within a few thousand iterations for both system sizes. However, achieving the next order of accuracy, 0.001, often requires 5–7 times more iterations, as the step size along the gradient direction becomes very small when the optimization approaches the local minimum.
Appendix D energy gap table
In Table 1, we provide the energy gap values for the two-dimensional $\sqrt{8}\times\sqrt{8}$ periodic cluster [41] for both the standard and disordered Hubbard Hamiltonians studied in work.
$S=0$ $S=1$ $S=2$ $S=3$
standard 1.12945 0.62746 0.78693 1.63385
disordered 0.56782 0.16508 0.32210 2.12290
Table 1: Energy gap values (in units of $\bar{t}$) for the standard and disordered Hubbard Hamiltonians for different total spin $S$ values.
For the standard Hubbard Hamiltonian, the total spin values corresponding to each energy gap are ranked from largest to smallest as follows: $S=3$, $S=0$, $S=2$, and $S=1$. This ordering corresponds to the observed difficulty of obtaining the ground state in each total spin $S$ sector, as seen in Fig. 3.
Similarly, in the disordered Hubbard Hamiltonian, we observe the same pattern, where the difficulty of finding the ground state is inversely related to the energy gap.

### v2

@@ IV Conclusion
We show that the simplification to an instantaneous interaction with a classical reservoir is an efficient method for ground-state preparation, offering a feasible approach for implementation in near-term quantum computers.
When comparing this work to previous approaches, we find that the classical reservoir method produces an ansatz expression analogous to the NP ansatz or the HVA ansatz [34, 33]. However, the motivations of the two approaches are different—our method arises from examining cooling dynamics while the HVA and NP ansatz are motivated by the adiabatic theorem [34, 33]. The classical reservoir method further enhances the efficiency of the algorithm by exploiting total spin symmetry and the closure property of the Lie algebra in the Hubbard hopping terms. Depending on hardware constraints, our algorithm can be implemented in a similar fashion to NP ansatz  [34, 36].
The classical reservoir method shows promise in preparing more accurate ground states in the strongly correlated regime, while simultaneously reducing implementation costs in the following ways:
@@
    The closure property of the Lie algebra allows us to use a one-dimensional indexing scheme, eliminating the need for fermionic swap (fSWAP) network [51, 48] to handle any hopping terms that require Jordan-Wigner strings in the ansatz for the state preparation.
Using the classical reservoir method, we find high-accuracy results —infidelities reach the range of $10^{-3}-10^{-6}$ for 8- and 10-site two-dimensional systems, even in the strong correlation region ($U/t{=}8$). We conjecture that this robustness arises from the fact that we do not start with the $U=0$ ground state as many other algorithms do, and thereby all $U$ values are essentially equal in terms of preparation difficulty.
V Acknowledgments
@@ Appendix A Initial state examples
The only requirement for the initial state is to have the target value of total spin. But, we find it is advantageous to intersperse the doubly occupied sites and the empty sites, so that the hopping terms in the ansatz can reduce the double occupancy in the first layer. Hence, we select the simplest states that fulfill this requirement, which is a product state. We illustrate this with an eight-site example.
@@ Appendix B Numerical optimization details
Figure B1: Gradient norms in panels (a) and (b) and relative errors in panels (c) and (d) during optimization over different numbers of layers. The $x$-axis represents the iteration index. The left panels, (a) and (c), correspond to the $2\times 4$ system, while the right panels, (b) and (d), correspond to the $2\times 5$ system.
We use the L-BFGS algorithm from the GPU-supported PyTorch package for the optimization [53]. The choice of this optimizer is empirical, as L-BFGS, a second-order gradient method using an approximate Hessian matrix, often outperforms first-order optimizers such as ADAM. This preference has also been observed in many other variational studies [34, 23, 54, 55].
The parameter amplitudes are initialized as random values within the range $[-0.001,0.001]$. By choosing a small random initialization, we ensure that the final optimized energy exhibits only a small numerical variation. The precise variance in the final energy due to the initial guess is case-dependent, as factors such as the potential parameter $U$ and the number of sites can contribute to this variation. In this work, using the mentioned initial guess, we find that the variation in the final energy is typically at most on the order of $10^{-2}$.
Though the goal of this work is to implement this algorithm on a quantum computer, we are currently using classical computer simulations. For the exponential of the hopping terms, we use an SU(2) identity [56] to implement the hopping terms instead of directly calculating the matrix exponential, which speeds up the computation. This is given by
@@
In Fig. B1, we provide details of the convergence behavior for the two-leg ladder system with periodic boundary conditions, along with the number of iterations required as the system size increases. We present results for $U/t=8$ only, as the total spin value of the ground state differs in the $U/t=2$ regime for the $2\times 4$ and $2\times 5$ systems, which can affect performance. The results show that the optimization typically reduces the relative error to below 0.01 within a few thousand iterations for both system sizes. However, achieving the next order of accuracy, 0.001, often requires 5–7 times more iterations, as both the gradient and the step size along the gradient direction become very small when the optimization approaches the local minimum.